# DrivAerNet: A Parametric Car Dataset for Data-Driven Aerodynamic Design and Graph-Based Drag Prediction Аннотация

### v1

@@
[…]
DrivAerNet: A Parametric Car Dataset for Data-Driven Aerodynamic Design and Graph-Based Drag Prediction
Аннотация
This study introduces DrivAerNet, a large-scale high-fidelity CFD dataset of 3D industry-standard car shapes, and RegDGCNN, a dynamic graph convolutional neural network model, both aimed at aerodynamic car design through machine learning. DrivAerNet, with its 4000 detailed 3D car meshes using 0.5 million surface mesh faces and comprehensive aerodynamic performance data comprising of full 3D pressure, velocity fields, and wall-shear stresses, addresses the critical need for extensive datasets to train deep learning models in engineering applications. It is 60% larger than the previously available largest public dataset of cars, and is the only open-source dataset that also models wheels and underbody. RegDGCNN leverages this large-scale dataset to provide high-precision drag estimates directly from 3D meshes, bypassing traditional limitations such as the need for 2D image rendering or Signed Distance Fields (SDF). By enabling fast drag estimation in seconds, RegDGCNN facilitates rapid aerodynamic assessments, offering a substantial leap towards integrating data-driven methods in automotive design. Together, DrivAerNet and RegDGCNN promise to accelerate the car design process and contribute to the development of more efficient vehicles. To lay the groundwork for future innovations in the field, the dataset and code used in our study are publicly accessible at https://github.com/Mohamedelrefaie/DrivAerNet111Data will be uploaded on paper acceptance.
keywords:
Parametric Design, Computational Fluid Dynamics, Car Aerodynamics, Surrogate Modeling, Graph Neural Networks| Dataset | Size | Aerodynamics Data | Wheels/Underbody | Parametric | Design Parameters | Open-source | |||
| u | Modeling | ||||||||
| Baque el al. 2018 [7] | 2,000 | ✔ | ✗ | ✗ | ✗ | ✗ | ✔ | 21 (3D) | ✗ |
| Umetani et. al 2018 [40] | 889 | ✔ | ✗ | ✔ | ✔ | ✗ | ✗ | - | ✔ |
| Gunpinar et. al 2019 [16] | 1,000 | ✔ | ✗ | ✗ | ✗ | ✗ | ✔ | 21 (2D) | ✗ |
| Remelli et. al 2020 [28] | 1,400 | ✗ | ✗ | ✗ | ✔ | ✗ | ✗ | - | ✗ |
| Jacob et. al 2021 [19] | 1,000 | ✔ | ✔ | ✔ | ✗ | ✔ | ✔ | 15 (3D) | ✗ |
| Usama et. al 2021 [41] | 500 | ✔ | ✗ | ✗ | ✗ | ✗ | ✔ | 40 (2D) | ✗ |
| Rios et. al 2021 [31] | 600 | ✔ | ✔ | ✗ | ✗ | ✗ | ✗ | - | ✗ |
| Song et. al 2023 [36] | 2,474 | ✔ | ✗ | ✗ | ✗ | ✗ | ✗ | - | ✔ |
| Li et. al 2023 [22] | 551 | ✔ | ✗ | ✗ | ✔ | ✗ | ✔ | 6 (3D) | ✗ |
| 611 | ✔ | ✗ | ✗ | ✔ | ✗ | ✗ | - | ✗ | |
| Trinh et. al 2024 [39] | 1,121 | ✗ | ✗ | ✔ | ✔ | ✗ | ✗ | - | ✗ |
| DrivAerNet (Ours) | 4,000 | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | 50 (3D) | ✔ |
@@ 1 Introduction
Reducing fuel consumption and CO2 emissions through advanced aerodynamic design is crucial to the automobile industry. This can help with faster transition towards electric cars, complementing the 2035 ban on internal combustion engine cars and aligning with the ambitious goal of achieving carbon neutrality by 2050 to combat global warming [25, 9, 23]. In aerodynamic design, navigating through intricate design choices involves a detailed examination of aerodynamic performance and design constraints, which is often slowed down by the time-consuming nature of high-fidelity CFD simulations and experimental wind tunnel tests. High-fidelity CFD simulations can take days to weeks per design [6], while wind tunnel testing, despite its accuracy, is limited to examining only a handful of designs due to time and cost constraints. Data-driven approaches can alleviate this bottleneck by leveraging existing datasets to navigate through design and performance spaces, thereby speeding up the design exploration process and efficiently assessing aerodynamic designs.
Although recent advancements in data-driven approaches for aerodynamic design are promising, these typically concentrate on simpler 2D cases [14, 38] or lower-fidelity CFD simulations [36, 22, 7] and overlook the complexities inherent in real-world 3D designs and the challenges posed by high-fidelity CFD simulations. According to [17], simplifying car designs by excluding components like wheels and mirrors, and not modeling the underbody, leads to a significant underestimation of aerodynamic drag. Taking these elements into account increased the drag by more than 1.4 times, highlighting the importance of detailed modeling for accurate aerodynamic analysis. Additionally, there is a dearth of publicly available high-fidelity car simulation datasets, potentially slowing down research in data-driven method development, as each researcher may need extensive computational resources to create their own data and test their methods on them.
@@
To demonstrate the importance of large-scale datasets, we also develop a surrogate model for aerodynamic drag prediction based on Dynamic Graph Convolutional Neural Networks [42]. Our model, RegDGCNN, operates directly on extremely large 3D meshes, eliminating the necessity for 2D image rendering or Signed Distance Fields (SDF) generation. RegDGCNN’s ability to swiftly identify aerodynamic improvements opens new avenues for creating more efficient vehicles by streamlining the evaluation of design adjustments. It marks a significant step towards optimizing car designs more efficiently.
@@
The introduction of a surrogate model, named RegDGCNN, based on Dynamic Graph Convolutional Neural Networks for prediction of aerodynamic drag. RegDGCNN outperforms state-of-the-art attention-based models [3, 36] for drag prediction by 3.57 on ShapeNet benchmark dataset while using 1000 fewer parameters, and achieves an score of 0.9 on the DrivAerNet dataset.
@@
The structure of our paper is as follows: Section 2 provides an overview of related work. Section 3 presents the DrivAerNet dataset, detailing the numerical simulation methods, CFD results, geometric feasibility analysis, and dataset characteristics. In Section 4, we introduce our RegDGCNN approach using the Dynamic Graph Convolutional Neural Network for regression tasks. Section 5 examines the application of the RegDGCNN model for surrogate modeling of aerodynamic drag, comparing its performance on the DrivAerNet and ShapeNet datasets and underscoring the benefits of scaling the training dataset. The paper highlights the limitations of our study and suggests avenues for future research, followed by a conclusion that summarizes the results and key findings.
@@ 2 Related Work
This section starts with an overview of aerodynamics datasets, and then transitions to discussing recent progress in 3D learning for aerodynamics.
@@ 2.1 Aerodynamics Datasets
Data-driven aerodynamic design is a methodology that leverages computational models and machine learning algorithms to optimize car shapes based on large volumes of aerodynamic performance data, aiming to improve efficiency and performance. A common type of data-driven aerodynamic design is surrogate modeling, which uses simplified models to approximate the behavior of complex aerodynamic phenomena, enabling faster simulations and iterations in the design process. It is particularly useful for preliminary design phases where quick evaluations are necessary. However, a significant portion of existing research on data-driven aerodynamic design is concentrated on simplified 2D scenarios such as airfoils/2D geometries [8, 38, 14, 41, 20, 16] or simplified 3D models [7, 28, 22, 40, 31, 36, 39]. While these studies are instrumental in understanding the fundamental physics, they often fall short in terms of applicability to complex 3D real-world problems. This gap is further widened by the absence of proprietary data from the industry, posing challenges in replicating and validating research findings. Moreover, the absence of a standardized benchmark dataset in the field hampers the ability to consistently evaluate and compare the efficacy of various machine learning methodologies. This stands in contrast to fields such as image processing or 3D shape analysis, where benchmark datasets like ImageNet [13] or ShapeNet [10] have accelerated significant advancements in deep learning by providing a common ground for methods comparison. Although benchmarks for CFD simulations and turbulence modeling [4, 5, 6] do exist, their limited size often renders them inadequate for the demands of deep learning techniques, underscoring the necessity for a broader and more comprehensive dataset.
Table 1 provides an overview of existing datasets in the literature for data-driven aerodynamic design. It compares their size, the inclusion of aerodynamic information such as drag coefficient (), lift coefficient (), velocity field (u), pressure (), whether they are parametric, the number of design parameters, and their availability as open-source. In addition, we consider the modeling of the rotating wheels and underbody for evaluation. Our dataset stands out with the largest size of 4,000 samples, comprehensive aerodynamic information, parametric details, modeling of the wheels and underbody, and open-source accessibility.
@@
2.2 Advancements in 3D Learning for Aerodynamics
The study from [39] presented a super-resolution model aimed at refining the estimated, yet coarsely resolved, flow fields around vehicles from deep learning predictions to a higher resolution, crucial for aerodynamic vehicle design. By incorporating a residual-in-residual dense block (RRDB) within the generator structure and employing a relativistic discriminator for enhanced detail capture, coupled with a novel distance-weighted loss and physics-informed loss to ensure physical accuracy, the methodology demonstrated a marked improvement in flow field enhancement around vehicles, outperforming previous approaches in this domain. Jacob et. al [19] demonstrated that deep learning, particularly a modified U-Net architecture using SDF, can accurately predict drag coefficients for a specific car design without the need for explicit parameterization.
In recent advancements in computational design, many studies have leveraged the ShapeNet [10] dataset for shape optimization and surrogate modeling in aerodynamics. ShapeNet is a dataset consisting of millions of 3D models spanning 55 common object categories, designed to support research in computer vision, robotics, and geometric deep learning. Using ShapeNet, [28] enhanced shape optimization by employing MeshSDF, offering a more flexible alternative to traditional hand-crafted parameterizations. Song et. al [36] introduced a novel 2D representation for 3D shapes, coupled with a surrogate model for aerodynamic drag, showcasing the potential for AI-driven design optimizations. Rios et. al [31] compared various design representation methods, including PCA, kernel-PCA, and a 3D point cloud autoencoder, highlighting the autoencoder’s capability for localized shape modifications in aerodynamic optimizations. The fourth study introduced the geometry-informed neural operator (GINO) [22], leveraging SDF, point clouds, and neural operators for efficient large-scale simulations, achieving significant speed-ups and error rate reductions in predicting surface pressure and drag coefficients on different car geometries.
@@ 2.2 Advancements in 3D Learning for Aerodynamics
Despite the novel methodologies employed in these studies, they faced limitations stemming from the inherent drawbacks of the ShapeNet dataset, such as lower mesh resolution, lack of watertight geometries, small dataset size, and oversimplifications like modeling cars as single-bodied entities without detailed considerations for components like wheels, underbody, and side mirrors, which can significantly impact real-world aerodynamic performance. This oversimplification can significantly impact real-world aerodynamic performance; according to [17], including these details in the DrivAer fastback model, increased the drag value from 0.115 to 0.278 in CFD simulations and from 0.125 to 0.275 in wind tunnel experiments. These increases represent a substantial increase in drag by approximately 142 and 120 respectively, underscoring the critical role of comprehensive modeling in achieving accurate aerodynamic assessments. Another common hurdle in both surrogate modeling and design optimization is the scarcity of data, which complicates efforts to replicate results or benchmark various models and approaches. Addressing this challenge, our contribution introduces DrivAerNet, a comprehensive benchmark dataset tailored for data-driven aerodynamic design, aiming to facilitate comparison and validation of future methodologies.

### v2

# DrivAerNet: A Parametric Car Dataset for Data-Driven Aerodynamic Design and Prediction DrivAerNet: A Parametric Car Dataset for Data-Driven Aerodynamic Design and Prediction

@@
Abstract
This study introduces DrivAerNet, a large-scale high-fidelity CFD dataset of 3D industry-standard car shapes, and RegDGCNN, a dynamic graph convolutional neural network model, both aimed at aerodynamic car design through machine learning. DrivAerNet, with its 4000 detailed 3D car meshes using 0.5 million surface mesh faces and comprehensive aerodynamic performance data comprising of full 3D pressure, velocity fields, and wall-shear stresses, addresses the critical need for extensive datasets to train deep learning models in engineering applications. It is 60% larger than the previously available largest public dataset of cars, and is the only open-source dataset that also models wheels and underbody. RegDGCNN leverages this large-scale dataset to provide high-precision drag estimates directly from 3D meshes, bypassing traditional limitations such as the need for 2D image rendering or Signed Distance Fields (SDF). By enabling fast drag estimation in seconds, RegDGCNN facilitates rapid aerodynamic assessments, offering a substantial leap towards integrating data-driven methods in automotive design. Together, DrivAerNet and RegDGCNN promise to accelerate the car design process and contribute to the development of more efficient cars. To lay the groundwork for future innovations in the field, the dataset and code used in our study are publicly accessible at https://github.com/Mohamedelrefaie/DrivAerNet.
Dataset | Size | Aerodynamics Data | Wheels/Underbody | Parametric | Design Parameters | Simulation Type | Mesh Resolution | Open-source | |||
u | Modeling | ||||||||||
Baque et al. 2018 [8] | 2,000 | ✔ | ✗ | ✗ | ✗ | ✗ | ✔ | 21 (3D) | RANS | - | ✗ |
Umetani et al. 2018 [48] | 889 | ✔ | ✗ | ✔ | ✔ | ✗ | ✗ | - | RANS | 700k | ✔ |
Gunpinar et al. 2019 [21] | 1,000 | ✔ | ✗ | ✗ | ✗ | ✗ | ✔ | 21 (2D) | RANS | 130k | ✗ |
Remelli et al. 2020 [36] | 1,400 | ✗ | ✗ | ✗ | ✔ | ✗ | ✗ | - | - | - | ✗ |
Jacob et al. 2021 [24] | 1,000 | ✔ | ✔ | ✔ | ✗ | ✔ | ✔ | 15 (3D) | LBM | 150M | ✗ |
Usama et al. 2021 [49] | 500 | ✔ | ✗ | ✗ | ✗ | ✗ | ✔ | 40 (2D) | - | - | ✗ |
Rios et al. 2021 [39] | 600 | ✔ | ✔ | ✗ | ✗ | ✗ | ✗ | - | RANS | 2M | ✗ |
Song et al. 2023 [44] | 2,474 | ✔ | ✗ | ✗ | ✗ | ✗ | ✗ | - | URANS | 3M | ✔ |
Li et al. 2023 [28] | 551 | ✔ | ✗ | ✗ | ✔ | ✗ | ✔ | 6 (3D) | RANS | 7.2M | ✗ |
611 | ✔ | ✗ | ✗ | ✔ | ✗ | ✗ | - | RANS | 600-700k | ✗ | |
Trinh et al. 2024 [47] | 1,121 | ✗ | ✗ | ✔ | ✔ | ✗ | ✗ | - | - | - | ✗ |
DrivAerNet (Ours) | 4,000 | ✔ | ✔ | ✔ | ✔ | ✔ | ✔ | 50 (3D) | RANS | 8-16M | ✔ |
@@ 1 Introduction
Reducing fuel consumption and CO2 emissions through advanced aerodynamic design is crucial for the automobile industry. This can facilitate a faster transition towards electric cars, complementing the 2035 ban on internal combustion engine cars and aligning with the ambitious goal of achieving carbon neutrality by 2050 to combat global warming [31, 10, 29]. In aerodynamic design, navigating through intricate design choices involves a detailed examination of aerodynamic performance and design constraints, which is often slowed down by the time-consuming nature of high-fidelity Computational Fluid Dynamics (CFD) simulations and experimental wind tunnel tests. High-fidelity CFD simulations can take days to weeks per design [6], while wind tunnel testing, despite its accuracy, is limited to examining only a handful of designs due to time and cost constraints. Data-driven approaches can alleviate this bottleneck by leveraging existing datasets to navigate through design and performance spaces, thereby speeding up the design exploration process and efficiently assessing aerodynamic designs.
Although recent advancements in data-driven approaches for aerodynamic design are promising, these advancements typically concentrate on simpler 2D cases [17, 46, 7] or lower-fidelity CFD simulations [44, 28, 8] and overlook the complexities inherent in real-world 3D designs and the challenges posed by high-fidelity CFD simulations. According to [22], simplifying car designs by excluding components like wheels and mirrors, and not modeling the underbody, leads to a significant underestimation of aerodynamic drag. Taking these elements into account increased the drag by more than 1.4 times, highlighting the importance of detailed modeling for accurate aerodynamic analysis. Additionally, there is a dearth of publicly available high-fidelity car simulation datasets, potentially slowing down research in data-driven method development, as each researcher may need extensive computational resources to create their own data and test their methods on them.
@@
To demonstrate the importance of large-scale datasets, we also develop a surrogate model for aerodynamic drag prediction based on Dynamic Graph Convolutional Neural Networks [50]. Our model, RegDGCNN, operates directly on extremely large 3D meshes, eliminating the necessity for 2D image rendering or Signed Distance Fields (SDF) generation. RegDGCNN’s ability to swiftly identify aerodynamic improvements opens new avenues for creating more efficient cars by streamlining the evaluation of design adjustments. It marks a significant step towards optimizing car designs more efficiently.
@@
The introduction of a surrogate model, named RegDGCNN, based on Dynamic Graph Convolutional Neural Networks for prediction of aerodynamic drag. RegDGCNN outperforms state-of-the-art attention-based models [3, 44] for drag prediction by 3.57 on the ShapeNet benchmark dataset while using 1000 fewer parameters, and achieves an score of 0.9 on the DrivAerNet dataset.
@@
The structure of our paper is as follows: Section 2 provides an overview of related work. Section 3 introduces the fundamentals of turbulence modeling, beginning with the Navier-Stokes equations, Reynolds averaging, and the SST turbulence model, which is used for generating our dataset. Section 4 presents the DrivAerNet dataset, detailing the numerical simulation methods, CFD results, geometric feasibility analysis, and dataset characteristics. In Section 5, we introduce our RegDGCNN approach using the Dynamic Graph Convolutional Neural Network for regression tasks. Section 6 examines the application of the RegDGCNN model for surrogate modeling of aerodynamic drag, comparing its performance on the DrivAerNet and ShapeNet datasets and underscoring the benefits of scaling the training dataset. The paper highlights the limitations of our study and suggests avenues for future research, followed by a conclusion that summarizes the results and key findings.
@@ 2 Related Work
This section starts with an overview of aerodynamics datasets and then transitions to discussing recent progress in 3D learning for aerodynamics.
@@ 2.1 Aerodynamics Datasets
Data-driven aerodynamic design is a methodology that leverages computational models and machine learning algorithms to optimize car shapes based on large volumes of aerodynamic performance data, aiming to improve efficiency. A common type of data-driven aerodynamic design is surrogate modeling, which uses simplified models to approximate the behavior of complex aerodynamic phenomena, enabling faster simulations and iterations in the design process. It is particularly useful for preliminary design phases where quick evaluations are necessary. However, a significant portion of existing research on data-driven aerodynamic design is concentrated on simplified 2D scenarios such as airfoils/2D geometries [9, 46, 17, 49, 25, 7, 21] or simplified 3D models [8, 36, 28, 48, 39, 44, 47]. While these studies are instrumental in understanding the fundamental physics, they often fall short in terms of applicability to complex 3D real-world problems. This gap is further widened by the absence of proprietary data from the industry, posing challenges in replicating and validating research findings. Moreover, the absence of a standardized benchmark dataset in the field hampers the ability to consistently evaluate and compare the efficacy of various machine learning methodologies. This stands in contrast to fields such as image processing or 3D shape analysis, where benchmark datasets like ImageNet [15] or ShapeNet [12] have accelerated significant advancements in deep learning by providing a common ground for method comparison. Although benchmarks for CFD simulations and turbulence modeling [4, 5, 6] do exist, their limited size often renders them inadequate for the demands of deep learning techniques, underscoring the necessity for a broader and more comprehensive dataset.
Table 1 provides an overview of existing datasets in the literature for data-driven aerodynamic design. It compares their size, the inclusion of aerodynamic information such as drag coefficient (), lift coefficient (), velocity field (u), pressure (), whether they are parametric, the number of design parameters, and their availability as open-source. In addition, we consider the modeling of rotating wheels and underbody for evaluation. The table also introduces details on the type of simulation employed, distinguishing between Reynolds-Averaged Navier-Stokes (RANS), Unsteady Reynolds-Averaged Navier-Stokes (URANS), and Lattice Boltzmann Method (LBM), as well as the mesh resolution used in each dataset. This comprehensive comparison highlights the strengths and limitations of each dataset in terms of aerodynamic performance and computational resources. Our dataset stands out with the largest size of 4,000 samples, comprehensive aerodynamic information, parametric details, modeling of the wheels and underbody, and open-source accessibility.
@@
2.2 Advancements in 3D Learning for Car Aerodynamics
The study from [47] presented a super-resolution model aimed at refining the estimated, yet coarsely resolved, flow fields around cars from deep learning predictions to a higher resolution, crucial for aerodynamic car design. By incorporating a residual-in-residual dense block (RRDB) within the generator structure and employing a relativistic discriminator for enhanced detail capture, coupled with a novel distance-weighted loss and physics-informed loss to ensure physical accuracy, the methodology demonstrated a marked improvement in flow field enhancement around cars, outperforming previous approaches in this domain. Jacob et al. [24] demonstrated that deep learning, particularly a modified U-Net architecture using SDF, can accurately predict drag coefficients for a specific car design without the need for explicit parameterization. Kontou et al. [27] introduced a DNN-based surrogate model for turbulence and transition closure in RANS equations, optimized using a Metamodel-Assisted Evolutionary Algorithm (MAEA). The DNN replaces the numerical solution of the turbulence model’s partial differential equations (PDEs), significantly reducing computational time in CFD-based shape optimization tasks. The RANS-DNN model demonstrated notable performance improvements in 3D aerodynamic cases, including the DrivAer car model [23], achieving efficient and accurate predictions of pressure fields and optimized geometries.
In recent developments in computational design, many studies have leveraged the ShapeNet [12] dataset for shape optimization and surrogate modeling in aerodynamics. ShapeNet is a dataset consisting of millions of 3D models spanning 55 common object categories, designed to support research in computer vision, robotics, and geometric deep learning. Using ShapeNet, Remelli et al. [36] enhanced shape optimization by employing MeshSDF, offering a more flexible alternative to traditional hand-crafted parameterizations. Song et al. [44] introduced a novel 2D representation for 3D shapes, coupled with a surrogate model for aerodynamic drag prediction, showcasing the potential for AI-driven design optimizations. Rios et al. [39] compared various design representation methods, including PCA, kernel-PCA, and a 3D point cloud autoencoder, highlighting the autoencoder’s capability for localized shape modifications in aerodynamic optimizations. The geometry-informed neural operator (GINO) [28] was introduced, leveraging SDF, point clouds, and neural operators for efficient large-scale simulations, achieving significant speed-ups and error rate reductions in predicting surface fields (e.g., pressure and wall shear stress) and drag coefficients of different car geometries. Deng et al. [16] proposed 3D Geometry-guided conditional adaptation (3D-GeoCA), a model-agnostic framework that enhances surrogate models for solving PDEs on arbitrary 3D geometries. Applied to the ShapeNet Car [48] and Ahmed-Body [28] datasets, 3D-GeoCA significantly reduces the L-2 error in large-scale Reynolds-Averaged Navier-Stokes simulations, offering a robust solution for handling complex geometries in aerodynamic modeling.
Building on these developments, transformers have recently been applied to PDEs with the introduction of Transolver [53], a novel approach utilizing Physics-Attention to group mesh points based on similar physical states. This method enables more efficient and accurate modeling of complex geometries, achieving state-of-the-art performance on the ShapeNet dataset [48].
@@ 2.2 Advancements in 3D Learning for Car Aerodynamics
Despite the novel methodologies employed in these studies, they faced limitations stemming from the inherent drawbacks of the ShapeNet dataset, such as lower mesh resolution, lack of watertight geometries, small dataset size, and oversimplifications like modeling cars as single-bodied entities without detailed considerations for components like wheels, underbody, and side mirrors, which can significantly impact real-world aerodynamic performance. According to Heft et al. [22], including these details in the DrivAer fastback model increased the drag value from 0.115 to 0.278 in CFD simulations and from 0.125 to 0.275 in wind tunnel experiments. This represents a substantial increase in drag by approximately 142 and 120 respectively, underscoring the critical role of comprehensive modeling in achieving accurate aerodynamic assessments. Another common hurdle in both surrogate modeling and design optimization is the scarcity of data, which complicates efforts to replicate results or benchmark various models and approaches. Addressing this challenge, our contribution introduces DrivAerNet, a comprehensive benchmark dataset tailored for data-driven aerodynamic design, aiming to facilitate comparison and validation of future methodologies.
3 Turbulence Modeling
In this section, we start with the basics of turbulence modeling, beginning with the Navier-Stokes equations. Next, we discuss Reynolds averaging, a key method in turbulence modeling. We then introduce the SST turbulence model, chosen for its effectiveness in accurately simulating a wide range of conditions, which makes it ideal for creating our dataset. The following derivations are summarized from [52, 30, 45].
3.1 Reynolds-Averaged Navier–Stokes Equations
The Navier-Stokes equations model fluid flow across various applications, from laminar to turbulent motion in science and engineering. Based on Newton’s second law, they describe fluid dynamics but are highly nonlinear and coupled, making analytical solutions infeasible except in simple cases. Numerical methods and CFD techniques are typically required for solving these second-order PDEs. The component-wise form of the Navier-Stokes equations for incompressible flow in Cartesian coordinates is essential for practical applications, allowing for the detailed modeling of fluid behavior in three-dimensional space.
Here, represent the velocity components in the directions, respectively, while is the pressure, is the fluid density, is the dynamic viscosity, and are the body force components acting in the respective directions. The terms account for the fluid’s inertia, pressure gradient, viscous diffusion, and external forces. The mass conservation and impulse conservation can be written in Einstein summation notation as:
Reynolds-Averaging
Reynolds averaging separates the fluctuating and mean components of a turbulent flow, simplifying analysis by focusing on the mean flow properties. It is widely used in turbulence modeling for velocity, pressure, and other flow quantities. The incompressible Navier-Stokes equations for the mean values of the velocity field under Reynolds averaging are given by:
This equation represents the conservation of momentum for the mean velocity field , including contributions from the Reynolds stresses , which are defined as the average of the product of velocity fluctuations and .
The Reynolds stress tensor plays a critical role in these equations, capturing the effects of turbulence on the mean flow.
This equation represents the incompressibility condition, ensuring that the divergence of the mean velocity field is zero.
Shear Stress Transport (SST)
The Shear Stress Transport (SST) model is a two-equation turbulence model that focuses on turbulence kinetic energy () and specific dissipation rate (). It is designed to address the limitations of the standard model, particularly its sensitivity to freestream values of and . The SST model is particularly effective at accurately depicting flow separation phenomena, making it a popular choice in aerodynamic and hydrodynamic simulations.
Model Equations
The turbulence specific dissipation rate equation is given by:
where:
is the diffusivity of .
represents the generation of turbulence kinetic energy due to mean velocity gradients.
and are model constants.
is the source term for .
The equation for turbulence kinetic energy is given by:
where:
is the diffusivity of .
is a model constant.
is the source term for .
The turbulence viscosity is obtained using the following relation:
where:
and are model constants.
is a blending function that ensures the model transitions smoothly between the model behavior in the near-wall region and the model behavior in the free stream.
represents the magnitude of the mean rate-of-strain tensor.
The SST model integrates distinct advantages in fluid dynamics simulations by employing a blending function to effectively merge the strengths of the model, which excels in near-wall regions, with the model, preferred for free-stream and far-field applications. This model stands out for its precise predictions of flow separation under adverse pressure gradients, enhancing its utility across various engineering disciplines including aerodynamics and hydrodynamics [52, 30].
For the DrivAerNet dataset generation, the SST turbulence model will be used to conduct the high-fidelity CFD simulations.